Chronium (II) sulfate, CrSO4, is a reagent that has been used in certain applications to help reduce carbon-carbon double bonds (c=c) in molecules to single bonds (c-c). The reagent can be prepared by the following reaction.


Zn(s) + K2Cr2O7(aq) + H2SO4(aq) → ZnSO4(aq) + CrSO4(aq) + K2SO4(aq) + H2O


What is the maximum number of grams of CrSO4 that can be made from a reaction mixture containing 3.2 mol Zn, 1.7 mol K2Cr2O7 and 5.0 mol H2SO4? 


Show solutions.


1
Expert's answer
2021-05-05T01:41:43-0400

4Zn(s) + K2Cr2O7(aq) + 7H2SO4(aq) → 4ZnSO4(aq) + 2CrSO4(aq) + K2SO4(aq) + 7H2O

M (CrSO4) = 148.1 g/mol

n = m / M

m = n x M

m1Zn(CrSO4) = 3.2 / 4 x 2 x 148.1 = 237 g

m1K2Cr2O7(CrSO4) = 1.7 x 2 x 148.1 = 503.5 g

m1H2SO4(CrSO4) = 5 / 7 x 2 x 148.1 = 211.6 g

H2SO4 is the limiting reactant for this reaction. The maximum number of grams of CrSO4 that can be made from a reaction mixture is 211.6 g.